Minutes of April 13, 2005 meeting

Andy Martinez, Megan Ingegno, Mary Brodman, Irina Loutchkina, Robert Krack, Matt Spaventa, Edward Suarez, Penny Weniger, Jeff Teichmann, Eugene McElroy, Martin Katz, Melanie Robinson, Krista Dandurand, Nicole Warren, Paul Young, Stacey DeMatteo, Shirley Peck, Melody Tomaszewicz, Brian Stubbs
Roselyn Riley



Brian/Alex- Busy with shelf reading of stacks collection. Dealt with several leaks in the stacks due to heavy rainfall.

Eugene/SMLR- SMLR had a complete security gate installed by 3M. Gate was purchased from department which was had surplused theirs. Special thanks to Andy and Jeff for being instrumental in the acquisition and installation of the gate.

-Rob/Alex- The Goffen collection (a gift donation) has been unpacked and is ready to be assessed by the appropriate selectors. There was a leak from the ceiling that seeped underneath the shelving ranges, some books were damaged but have been saved. A patch of mold was found on a wall in Collection Services caused by a leak from a toilet on the floor above, the mold (and a section of the wall) were removed and Facilities is investigating to see if there is more mold. Repairs are underway.

-Matt/Douglass- The shelf reading of the collection has been slowed because of a shortage of student help. The East Room shift was completed over Spring Break. All of the stack range guides in the East Room have been revised using the new call numbers and the standard range guides available on the CMFG website.

-Eddie/LSM- All of the items from the Special Collections Room (X, X2, X-Per, X-Folio) have been transferred online to the Annex or SC/UA. The majority of the items have been shipped, except for two carts which need to be relabeled before they go to the Annex. The future plans for the room have yet to be decided. Busy assessing space on the 2nd floor created by the weeding project and putting hooks and ends where needed. Before pulling for weeding project resumes, all parties will meet to discuss how to proceed.

-Paul/NBL- Still regularly receiving DOCNJ at Alex as well as foreign and international items at LSM. A shift may be needed in the summer to accommodate additions. Paul will let the group know if assistance is needed.

-Penny/Dglss-Toured the Annex and reviewed their space management database, which was impressive. Penny’s past experience at Princeton ReCAP was beneficial in comparing Annex to that facility. The shelf space database is on pa with what other book storage facilities pay large sums of money to outside vendors for. Along with Jeff has been working on finalizing the PowerPoint’s for the Mold Awareness Program to be held at the end of May. Was able to review two facilities databases to collect data for the standard floor map project, but information was outdated or not useful so the map will need to be created from scratch.

-Stacy/Annex-Still receiving transfers from the LSM weeding project. Was able to complete the processing of the transfers from the LSM Special Collections Room Project.

-Irina/Alex- Working on the In-Process books, which have greatly increased in the past two months.

-Nicole/Chang- Completed an inventory and shelf reading of entire collection over spring break. Working off of inventory reports, a number of missing items were found.

-Jeff/Alex- Announced that Christopher Lee will be leaving NBL for a position at Seton Hall University. His last day is 4/22 and a farewell luncheon will be held for him 4/19 at the University Inn and Conference Center.

-Krista/KLMR- A large book sale will be held during the week of 4/18, including items withdrawn from KLMR and LSM weeding projects. The book sale will take place on the library patio, with most items going for $1.

-Melody/CS- In March/2005, 1757 books were processed from the LSM weeding project. 11,000 items have been processed since the project began. A small weeding project of the Reference Collection at the Math Library is in the planning stages. Progress continues on the Serials Inventory project.

-Mary/Alex-The shelf reading and mini-shifting on 2B will be completed soon. New range guides are being placed as the mini-shifts are completed. The standard range guide templates will not be used until the entire collection is shifted. The student appreciation celebration was a big success.

-Shirley/Alex- Has encountered an increase in stacks searches, missing book slips and claims returned, since the end of the semester is nearing.

-Melanie/Math- Completed a mold awareness handout, as part of the sub-group working on the Mold Awareness Program. Continues to work on the Serials Inventory Project.


-Eddie and Rob are in the process of revising and updating the content of the Library of Congress Classification System PowerPoint presentation. The PowerPoint to be revised resides on the Collection Management part of the NB Access Services website, under Documentation:

-Once revised, the PowerPoint will be an excellent tool for use by our part time employees.

-A draft will be made available for the group to review by our next meeting.


-The sub-group (members from the CMFG and FSFG) met on 3/31 to discuss the development of this project.

-Two types of floor plans, one to be available electronically and one to be made available as a handout in the units, were decided on.

-The electronic floor plan will make use of Flash and imbedded pictures to deliver information on collections and other points of interest in the libraries while the handouts will be to scale.

-The group will meet again in May. CMFG members are encouraged to forward their ideas and comments on standard floor maps to either Penny or Andy.

-Chang and Kilmer CM staff were asked to forward a copy of the existing floor maps to Andy. All available floor maps for the units will be made available for review by the group on the Alex t-drive:
(Alex t-drive/Common/NB Col Mgmt Group/Floor Plan Sub Group).


-The entire shift was completed within the specified time period during Spring Break/2005. A team of 3 shifters worked an hourly work schedule during the shift.

+16 double-sided ranges (some with 8 sections, some with 9 sections) were shifted to maximize space. Seven inches of growth space were left on the end of each shelf.

-Average time to shift one side of a range was 81 minutes. On average, more time was spent on shifting each range at the beginning of the project.

-As part of the shift, each shelf was cleaned with warm water, removing all of the dust buildup.

-Once the shift of all of the items in the room was complete, approximately 1 ˝ ranges of the Tapestry Room’s contents were moved into the East Room. The East Room now contains call numbers A – HC110.

-As part of the shift, the new standard range guide templates were used to revise the range guides in the room. Work still needs to be done on adjusting the templates so that cards for the entire collection can be saved, rather than four at a time.

-A digital picture of the standard stacks range guides that were installed at the Douglass Library was distributed for the group for review. This picture may be posted to the CMFG part of the NB Access Services page as an example of the guides in action.

-a shift of a portion of the West Room has been proposed to take place during the summer.

-As part of the shift, shelving was emptied in the Tapestry Room. Library Interiors was contracted to dismantle this shelving and move it to the Music Library. This project was completed on 4/8/05.


The reports for all RUL units will be available by Friday 4/15/05 for review. The reports cover the dates 7/1/04 to 3/15/05. All of the reports will be posted to the Alex t-drive:
(Alex t-drive/Common/MISSING BOOK REPORTS [Access Services}/)

A spreadsheet with the breakdown of each report was distributed to the group for review. The spreadsheet listed each unit, with the breakdown for the Lost-Paid, Lost-Replace, and Missing reports. The statistical summaries of all of the reports generated for the fiscal year are also available on the Alex t-drive.

One major change in the format of the reports is the separation of the Alexander missing report into two reports. One report will include all of the Alex missing items, excluding any recon items. The second report will only include the Alex missing recon items. CM staff at Alexander will search the missing recon list one final time, it will then be forwarded to the Annex for searching, and then Systems will remove all of the remaining missing recon items.

The next comprehensive RUL missing report will be the yearly report for withdrawal generated in July/2005.


Ian, Tim, and Andy have finished installing temperature/humidity readers in most of the RUL units. A reference sheet listing each unit, location of readers and date placed was completed and made available to the group for reference.

The group was reminded that constant temperature and humidity readings are available with the LCD readers. Any large fluctuations in temperature or humidity should be reported to Administration.

With these data loggers, a temperature and humidity reading is taken every couple of seconds and recorded. Ian plans on downloading this information once a month and creating graphs of the readings for analysis.

Andy reminded the group to inform student staff about the location of the readers and to try to keep an eye on the equipment. Try to incorporate spot checks of the equipment in the shelving schedules of your part time employees.

Please let Ian know about any water leaks that occur in the general area of one of the data loggers. An email is fine.


Wednesday May 11, 2005
Place to be announced

URL: http://www.libraries.rutgers.edu/rul/staff/groups/collection_mgt_group/minutes/collection_mgt_group_05_04_13.shtml
Website Feedback  |  Privacy Policy
© Copyright 1997-, Rutgers University Libraries